# Prague, Czech

### v2

@@ 2 Related Work
There exists a variety of theoretical rounding error analyses of basic mixed-precision algorithms aimed at understanding the underlying arithmetic behavior; see the survey doi:10.1177/10943420211003313 and references therein. However, for real-world problems, which can be more complex, the scientific community typically uses more general approaches such as precision tuning tools, which operate at program compile time and runtime, immediately resulting in reduced-precision programs produced automatically. In this section, we briefly review existing approaches for the analysis of mixed-precision algorithms and automated precision tuning that are closely related to our method.
Theoretical analysis aims to design reduced-precision settings for specific steps and analyze their rounding error so that the reduced-precision arithmetic can be used safely under certain assumptions to retain algorithm accuracy and numerical stability. Analyses in WILKINSON1966 and 10.1145/321386.321394 demonstrate that in standard iterative refinement, residuals computed in extra precision (typically twice the working precision) can guarantee the convergence of the relative error to the level of the working precision if the matrix $A$ has a condition number safely less than the reciprocal of the unit-roundoff, $u$. More recently, doi:10.1137/17M1122918 found that it is possible to produce solutions to iterative refinement with normwise relative error of order $u$ for systems with condition numbers of order $u^{-1}$ or larger by using a GMRES-based approach called GMRES-IR (see Algorithm 2). The authors extended this work in doi:10.1137/17M1140819 to develop a general analysis of iterative refinement using up to three difference precisions, as well as a general solver. This work shows that, depending on constraints on the condition number $\kappa(A)$, the most expensive part of the computation can often be carried out entirely in low precision without harming the accuracy. However, to obtain theoretical results such as these, a great amount of effort is needed, and the analysis often relies on knowledge about the problem, e.g., the condition number, which may not be readily available in practice. To circumvent these obstacles, one can switch to precision tuning tools that operate during compile time and runtime; see the survey 10.1145/3381039 and references therein.
@@ 4.2 Contextual Bandit Formulation for Precision selection
The reward $r(s,\mathbf{a})$ has the form

### v3

@@ 2 Related Work
There exists a variety of theoretical rounding error analysis of basic mixed-precision algorithms aimed at understanding the underlying arithmetic behavior; see the survey doi:10.1177/10943420211003313 and references therein. However, for real-world problems, which can be more complex, the scientific community typically uses more general approaches such as precision tuning tools, which operate at program compile time and runtime, immediately resulting in reduced-precision programs produced automatically. In this section, we briefly review existing approaches for the analysis of mixed-precision algorithms and automated precision tuning that are closely related to our method.
Theoretical analysis aims to design reduced-precision settings for specific steps and analyze their rounding error so that the reduced-precision arithmetic can be used safely under certain assumptions to retain algorithm accuracy and numerical stability. Analyses in WILKINSON1966 and 10.1145/321386.321394 demonstrate that in standard iterative refinement, residuals computed in extra precision (typically twice the working precision) can guarantee the convergence of the relative error to the level of the working precision if the matrix $A$ has a condition number safely less than the reciprocal of the unit-roundoff, $u$. More recently, doi:10.1137/17M1122918 found that it is possible to produce solutions to iterative refinement with normwise relative error of order $u$ for systems with condition numbers of order $u^{-1}$ or larger by using a GMRES-based approach called GMRES-IR (see Algorithm 2). The authors extended this work in doi:10.1137/17M1140819 to develop a general analysis of iterative refinement using up to three difference precisions, as well as a general solver. This work shows that, depending on constraints on the condition number $\kappa(A)$, the most expensive part of the computation can often be carried out entirely in low precision without harming the accuracy.
However, to obtain theoretical results such as these, a great amount of effort is needed, and the analysis often relies on knowledge about the problem, e.g., the condition number, which may not be readily available in practice. To circumvent these obstacles, one can switch to precision tuning tools that operate during compile time and runtime; see the survey 10.1145/3381039 and references therein.
@@ 4.2 Contextual Bandit Formulation for Precision selection
The reward $r(s,\mathbf{a})$ is designed to lead RL to choose precision configurations that lead the linear solvers to have an acceptable level of converged iterations as well as accuracy, it has the form:
@@ 5.3 Simulations on Sparse Systems
5.4 Ablation Study
To further verify the effectiveness of our RL modeling for precision selection. We focus on the converged total GMRES iterations taken for the iterative refinement and conduct an ablation study on training with (19) without $f_{\text{penalty}}$—the penalty on iteration terms.
Figure 3: Average Floating-point Types Selected Across Low, Medium, and High Condition Ranges.
The results are presented in Table 5.4, together with Figure 3 showing precision types chosen by RL. Similarly to above, we observed that between low- and medium-conditioned linear systems, RL strategically selects aggressive reduced-precision steps, whereas in high-conditioned linear systems it adopts a more conservative strategy for precision selection. The removal of it $f_{\text{penalty}}$ results in increasing iterations for the RL agent to achieve comparable accuracy, which means the iterative refinement uses extra iterations to compensate for the accuracy loss resulting from using increased low-precision steps. The results indicate the importance of the reward function modeling for RL training and demonstrate the effectiveness of RL for precision selection under proper modeling.
Table 7: Average Performance Metrics Across Condition Ranges for Dense Systems (Ablation study).
$\tau=10^{-6}$
RL($W_{1}$) Low ($10^{0}$$10^{3}$) 100 % $8.86\text{\times}{10}^{-08}$ $1.45\text{\times}{10}^{-10}$ $2.65$ $11.05$
Medium ($10^{3}$$10^{6}$) 100 % $4.22\text{\times}{10}^{-07}$ $4.97\text{\times}{10}^{-11}$ $2.54$ $5.94$
High ($10^{6}$$10^{9}$) 100 % $1.94\text{\times}{10}^{-09}$ $1.63\text{\times}{10}^{-14}$ $2.32$ $3.25$
RL($W_{2}$) Low ($10^{0}$$10^{3}$) 100 % $3.37\text{\times}{10}^{-07}$ $2.20\text{\times}{10}^{-08}$ $2.46$ $12.08$
Medium ($10^{3}$$10^{6}$) 100 % $4.22\text{\times}{10}^{-07}$ $4.97\text{\times}{10}^{-11}$ $2.54$ $5.94$
High ($10^{6}$$10^{9}$) 100 % $1.94\text{\times}{10}^{-09}$ $1.63\text{\times}{10}^{-14}$ $2.32$ $3.25$
[…]
$\tau=10^{-8}$
RL($W_{1}$) Low ($10^{0}$$10^{3}$) 100 % $1.34\text{\times}{10}^{-11}$ $1.62\text{\times}{10}^{-12}$ $2.38$ $15.97$
Medium ($10^{3}$$10^{6}$) 100 % $3.49\text{\times}{10}^{-06}$ $6.44\text{\times}{10}^{-11}$ $2.31$ $12.37$
RL($W_{2}$) Low ($10^{0}$$10^{3}$) 89.2 % $2.49\text{\times}{10}^{-07}$ $61\text{\times}{10}^{-08} 2.2$72.27 $21.03$
Medium ($10^{3}$$10^{6}$) 100 % $3.49\text{\times}{10}^{-06}$ $6.48\text{\times}{10}^{-11}$ $2.34$ $12.51$
High ($10^{6}$$10^{9}$) 100 % $3.91\text{\times}{10}^{-06}$ $4.28\text{\times}{10}^{-11}$ $2.18$ $4.29$
[…]